OpenStudy (anonymous):

If y = 8x − 2, which of the following sets represents possible inputs and outputs of the function represented as ordered pairs? {(0, −2), (1, 6), (2, 14)} {(0, −2), (2, 13), (−5, −40)} {(1, 7), (2, 7), (3, 7)} {(1, 6), (2, 13), (3, 24)}

OpenStudy (anonymous):

@DanJS @rhr12 @ali2x2

OpenStudy (danjs):

each of those pairs is a point (x,y) test all 4 points, they all have to satisfy y = 8x - 2

OpenStudy (anonymous):

so, whats the answer?

OpenStudy (danjs):

For example the first one... (0 , -2) -2 = 8*0 -2 true (1 , 6) 6 = 8*1 - 2 true (2 , 14) 14 = 8*2 - 2 true The first set of points works

OpenStudy (anonymous):

A?

OpenStudy (rhr12):

First one is the answer

OpenStudy (rhr12):

Yes

OpenStudy (danjs):

A works, not sure about the others if more

OpenStudy (rhr12):

Only a works. I did it.